OpenStudy (magbak):

I will award medal!!!!! Please Help!!!!!! Using complete sentences, describe how you would find all possible rational zeros of the polynomial function f(x) = 9x^4 – 17x^3 + 2x^2 – 3x + 33. Be sure to provide the answer in your explanation.

OpenStudy (magbak):

@whpalmer4

OpenStudy (anonymous):

ok so we are not going to worry about what the actual roots are, we are only going to answer the question that is asked

OpenStudy (anonymous):

the question is "Using complete sentences, describe how you would find all possible rational zeros of the polynomial function f(x) = 9x^4 – 17x^3 + 2x^2 – 3x + 33."

OpenStudy (magbak):

Yes but it asks for the answr I belive that means the roots

OpenStudy (magbak):

Yes no roots

OpenStudy (anonymous):

the possible rational roots are rational numbers, numbers that look like \(\frac{p}{q}\) were \(p\) and \(q\) are integers

OpenStudy (anonymous):

and \(p\) has to divide the constant, which in this case is \(33\) so your choices for \(p\) are \(\pm1,\pm3,\pm11,\pm33\)

OpenStudy (anonymous):

your choices for \(q\) are numbers that divide the leading coefficient, which in your case is \(9\) the choices for \(q\) are therefore \(\pm1,\pm3,\pm9\)

OpenStudy (anonymous):

make all possible combinations of those fractions

OpenStudy (anonymous):

if \(q=1\) they are \[\pm1,\pm3,\pm11,\pm33\]

OpenStudy (magbak):

Ok I will post a new question and tag you

OpenStudy (anonymous):

if \(q=3\) they are \[\pm\frac{1}{3},\pm\frac{11}{3}\] and some others that are already listed

OpenStudy (anonymous):

if \(q=9\) then you get a couple more, like \[\pm\frac{1}{9}\] and \[\pm\frac{11}{9}\]